OpenStudy (anonymous):

INTEGRAL HELP!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Will medal and make you a crown

OpenStudy (anonymous):

\[\int\limits_{2}^{x}\frac{ dt }{ t^2-1}\]

OpenStudy (anonymous):

Find the f(x) when \[\lim_{x \rightarrow \infty}\]

OpenStudy (anonymous):

I've been trying to solve it since yesterday

OpenStudy (anonymous):

Any thoughts fellas?

OpenStudy (anonymous):

@IrishBoy123 ?

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@Joel_the_boss

OpenStudy (anonymous):

any idea @Ziferot ?

OpenStudy (butterflydreamer):

hint: